OpenStudy (anonymous):

Simplify (x2 + 5x + 12)(x2 + 6x + 7).

OpenStudy (kidrah69):

Just combine the like terms :)

OpenStudy (kidrah69):

\[x2 + 5x + 12+x2 + 6x + 7\]

OpenStudy (anonymous):

can you explain how to do it?

OpenStudy (kidrah69):

All you do is look for like terms do you know what like terms are?

OpenStudy (anonymous):

x2 5x 12 x2 6x and7 ?

OpenStudy (kidrah69):

They are all terms

OpenStudy (anonymous):

oh then nope

OpenStudy (kidrah69):

We want LIKE terms so x^2 + x^2 6x+5x 7

OpenStudy (kidrah69):

now just add those :-)

OpenStudy (kidrah69):

What is x^2+x^2

OpenStudy (anonymous):

x^4

OpenStudy (kidrah69):

No

OpenStudy (kidrah69):

x=1 what is 1+1

OpenStudy (anonymous):

2

OpenStudy (kidrah69):

good so we have 2x^2 :) what is 6+5 ?

OpenStudy (anonymous):

11

OpenStudy (kidrah69):

Good so we now have a simplified statement! 2x^2+11x+7 :)

OpenStudy (anonymous):

but that isnt a answer choice for me

OpenStudy (kidrah69):

:/

OpenStudy (anonymous):

A. x4 + 37x2 + 84 B. x4 + 30x2 + 84 C. x4+ 11x3 + 49x2 + 107x + 84 D. x4 + 6x3 + 7x2 + 11x + 19

OpenStudy (kidrah69):

We were supposed to multiply

OpenStudy (kidrah69):

so x^2 times x^2=?

OpenStudy (kidrah69):

We need to foil everything :)

OpenStudy (anonymous):

4

OpenStudy (kidrah69):

x^4 good x^2 times 6x = ?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

6x^3?

OpenStudy (kidrah69):

yes thats right

OpenStudy (kidrah69):

x^2 times 7 =?

OpenStudy (anonymous):

7x^2

OpenStudy (kidrah69):

good :) so far we have x^4+6x^3+7x^2 do you see any answer choices that match up to this?

OpenStudy (er.mohd.amir):

x2(x2+6x+7)+5x(x2+6x+7)+12(x2+6x+7) then multply

OpenStudy (anonymous):

d.x4 + 6x3 + 7x2 + 11x + 19?

OpenStudy (er.mohd.amir):

gives x4+6x3+7x2+5x3+30x2+35x+12x2+72x+84

OpenStudy (anonymous):

but where do 11x and 19 come from?

OpenStudy (anonymous):



OpenStudy (anonymous):

by multiplying

OpenStudy (welshfella):

you use the distributive law first multiply all the terms in the second parentheses by x^2, then repeat using + 5x then finally by +12 Then combine like terms